OpenStudy (anonymous):

Find the slope and the y-intercept: 4x-3y=7 How do I solve this?

OpenStudy (yanasidlinskiy):

y=mx+b is the formula you need to find out the answer!:)

OpenStudy (anonymous):

isolate into slop intercept form, y=mx+b, so -3y=7-4x, then y= (7/-3)+(4/3)x so the y int is -(7/3) and the slope is 4/3

OpenStudy (anonymous):

How did you get the y= (7/-3)+(4/3)x portion?

OpenStudy (anonymous):

you divide the left side and right side by -3. the left side cancels out, leaving y, and the right side every coeff is divided by -3

OpenStudy (anonymous):

Okay, I understand that part now. How can I tell which one is the y intercept and which one is the slope?

OpenStudy (anonymous):

the mx is (4/3)x so the slope m, so (4/3). the y intercept is (7/-3) because it has no variable attached and it is a constant

OpenStudy (anonymous):

How do I make the y intercept into an ordered pair? Is it (7,-3)?

OpenStudy (anonymous):

no, the y intercept is always 0 for x. so it is (0, (7/-3)) it's just a fraction

OpenStudy (anonymous):

Oh! That makes so much sense now. Thank you so much, i'm going to keep practicing these problems. Thanks!! :D